### v1

# Ulrich complexity and Picard rank of bidouble Planes

Abstract.
We determine the Picard number and the Ulrich complexity of general bidouble covers of the projective plane, providing the first systematic study of Ulrich bundles on non-cyclic abelian covers. For a bidouble plane branched along three smooth curves of degrees $n_{1},n_{2},n_{3}$, we show that $\rho(S)=1$ unless $(n_{1},n_{2},n_{3})$ belongs to an explicit list, thereby extending Buium’s classical results on double planes to the non-cyclic case. As an application, we determine the range of branch degrees for which Ulrich line bundles could exist, and we show that every even bidouble plane carries a special rank two Ulrich bundle. Our method combines the invariant-theoretic decomposition of $H^{2}(S,{\mathbb{Q}})$ under the Galois group with cohomological criteria for Ulrich bundles.
Key words and phrases:
Picard number, Ulrich bundles, bidouble covers, singular double covers
2020 Mathematics Subject Classification:
Primary 14C22; Secondary 14C20, 14E20, 14J60
J.C. was supported by a Simons Foundation grant (Grant #550023). J.C-P. was supported by the Master’s Program, Universidad del Valle. S.T. was partially supported by the PRIN project “Multilinear Algebraic Geometry” of MUR (2022ZRRL4C)
1. Introduction
Ulrich bundles constitute a distinguished class of arithmetically Cohen–Macaulay vector bundles whose cohomological extremality reflects the embedding of a projective variety into projective space. Since their introduction by Eisenbud and Schreyer [15], Ulrich bundles have played a central role in algebraic geometry, appearing in the study of syzygies, Chow forms, and the geometry of linear projections (see Section 2.1 for the definition and basic facts).
Despite intensive progress, the existence and structure of Ulrich bundles on surfaces of general type are currently understood only in a limited number of cases. The case of cyclic covers of $\mathbb{P}^{2}$ has been particularly fruitful. Indeed, for double planes, Ulrich line bundles were studied in [22], while rank two Ulrich bundles were constructed in [25, 18]. More generally, existence results are now known for double coverings of projective spaces, and on Fano and Calabi–Yau double coverings of $\mathbb{P}^{3}$ [18, 27, 28] and for cyclic coverings of higher rank [23]. It is also worth mentioning the work of [3], where the authors investigate aCM sheaves on double planes, that is, hyperplanes of ${\mathbb{P}}^{3}$ with multiplicity two.
This paper provides the first systematic study of the Ulrich complexity of bidouble planes with respect to the pullback of a line. Although this divisor is not very ample, it is ample and globally generated, so the definition of Ulrich bundles extends naturally (see for example [7]). Bidouble planes are understood here as $(\mathbb{Z}/2\mathbb{Z})^{2}$-covers of $\mathbb{P}^{2}$ branched along three transversal curves whose degrees share the same parity. We call the bidouble plane even if all of these degrees are even; otherwise, we call it odd.
Our first main result determines the Picard number of a general bidouble plane, extending the classical results of Buium [6] from double planes to non-cyclic abelian covers.
Theorem 1.1.
Let $S\to\mathbb{P}^{2}$ be a smooth bidouble plane branched along general curves $D_{1},D_{2},D_{3}$ of degrees $n_{1},n_{2},n_{3}$, respectively. Then $\rho(S)>1$ if and only if some permutation of $(n_{1},n_{2},n_{3})$ belongs to
$\{(0,2,2n):n\geq 1\}\;\cup\;\{(0,4,2n):n\geq 2\}\;\cup\;\{(1,3,2n+1):n\geq 0\}\;\cup\;\{(2,2,2n):n\geq 1\}.$
The computation relies on the eigenspace decomposition of $H^{2}(S,\mathbb{Q})$ under the $(\mathbb{Z}/2)^{2}$-action and on a detailed analysis of the Hodge structure associated to the intermediate double covers determined by the three involutions of $S$. One consequence is that, for general choices of branch curves, the surface $S$ carries no Ulrich line bundles with respect to $\mathscr{O}_{S}(1)$: the Néron–Severi group is too small for such a bundle to exist (cf. Theorem 4.5). We therefore turn to the existence of rank two Ulrich bundles on $S$. The smallest integer $r\geq 1$ such that there exists a rank $r$ Ulrich bundle on $S$ with respect to $H$ is called the Ulrich complexity and is denoted by $\operatorname{uc}(S,H)$. Our second main theorem establishes the Ulrich complexity of even bidouble planes.
Theorem 1.2.
Let $S$ be a smooth bidouble plane of $\mathbb{P}^{2}$ branched along generic smooth curves of degrees $n_{1},n_{2},n_{3}$, and assume $n_{1}\leq n_{2}\leq n_{3}$. Define
$T_{1}=\{(0,4,2n)\mid n\geq 2\}\;\cup\;\{(2,2,2n)\mid n\geq 1\},\qquad\text{and}\qquad T_{2}=\{(0,2,2),\,(0,2,4)\}.$
If $S$ is an odd bidouble cover, then its Ulrich complexity with respect to $\mathscr{O}_{S}(1)$ satisfies $\operatorname{uc}(S,\mathscr{O}_{S}(1))>1$.
If $S$ is an even bidouble cover, then:
  • (a)
    If $(n_{1},n_{2},n_{3})\notin T_{1}\cup T_{2}$, then $\operatorname{uc}(S,\mathscr{O}_{S}(1))=2$.
  • (b)
    If $(n_{1},n_{2},n_{3})\in T_{1}$, then $\operatorname{uc}(S,\mathscr{O}_{S}(1))\leq 2$.
  • (c)
    If $(n_{1},n_{2},n_{3})\in T_{2}$, then $\operatorname{uc}(S,\mathscr{O}_{S}(1))=1$.
The construction of Ulrich rank-two vector bundles to prove Theorem 1.2 uses the Cayley–Bacharach–Serre correspondence method (cf. Theorem 5.1). Additionally, we compute the Ulrich complexity in some low-degree cases, completing the picture for even bidouble planes (cf. Proposition 4.6).
In [4, Theorem 1.1], Basu and Pal establish the existence of rank two Ulrich bundles on certain surfaces of general type. However, our results lie outside the scope of their framework, as they assume that the surfaces are polarized by their canonical bundle. This is not our situation, since we investigate Ulrich bundles with respect to a non–very ample divisor. Moreover, [4, Theorem 1.1] requires that the canonical class generate the Picard group.
The paper is organized as follows. In Section 2, we recall the basic facts about Ulrich bundles and bidouble covers. In Section 3, we study the Nerón–Severi group of bidouble planes: first, we show that the Picard number of a bidouble plane is determined by those of its intermediate covers; second, we establish the numerical conditions under which the Picard number is strictly larger than one, leading to the proof of Theorem 1.1. In Section 4, we show that the existence of an Ulrich line bundle forces the Picard number to be greater than one. For double planes, it was shown in [22] that the only obstruction is having Picard number one. In contrast, we exhibit families of bidouble covers with Picard number greater than one that nevertheless admit no Ulrich line bundles. On the positive side, we also construct explicit Ulrich line bundles in the cases $(n_{1},n_{2},n_{3})=(0,2,2)$ and $(0,2,4)$. Finally, in Section 5, we prove Theorem 1.2, showing that the Ulrich complexity of an even bidouble plane with sufficiently large degree is equal to $2$.
Remark 1.3.
We restrict to bidouble covers of $\mathbb{P}^{2}$ for the following reasons. Most of our non-existence results for Ulrich line bundles hinge on showing that the Picard number is one, which in turn forces the base surface to have a Picard number one. Although some arguments extend to more general bases, proving that the intermediate quotients also have Picard number one (cf. Lemma 3.4) requires realizing them as hypersurfaces in a weighted projective space. This situation arises precisely when the base is $\mathbb{P}^{2}$. Notice that one of our results (cf. Proposition 4.4) establishes the nonexistence of Ulrich line bundles on double covers of $\mathbb{P}^{1}\times\mathbb{P}^{1}$ branched along symmetric divisors, a fact that has not been previously recorded in the literature.
Acknowledgements
We thank Pedro Montero, Joan Pons-Llopis, and Roberto Villaflor for useful discussions. In addition, we thank Jorge Duque, Francesco Malaspina, Johannes Rau, David Rohrlich, and Giancarlo Urzúa for their comments on a previous version of this manuscript. The first author was supported by Simons Foundation grant no. 550023, the second author was partially supported by the Master’s Program, Universidad del Valle, and the third author was partially supported by the PRIN project “Multilinear Algebraic Geometry” of MUR (2022ZRRL4C). The third author also extends thanks to Rijksuniversiteit Groningen/Campus Fryslân for their hospitality.
2. Preliminaries
All varieties in this work are defined over the complex numbers $\mathbb{C}$.
2.1. Ulrich bundles
In recent years, considerable attention has been devoted to a class of vector bundles known as Ulrich bundles. For a smooth projective variety with an embedding $X\subset\mathbb{P}(|H|)$ given by a very ample divisor $H$, a vector bundle $\mathscr{E}$ on $X$ is called Ulrich with respect to $H$ if $H^{i}(X,\mathscr{E}(-jH))=0$ for all $i\geq 0$ and $1\leq j\leq\dim X$. We refer the reader to [12] for a comprehensive introduction to Ulrich bundles.
In this paper, we adopt a slightly more general perspective: rather than assuming that the polarization is very ample, we only require it to be ample and globally generated.
Definition 2.1.
Let $X$ be a smooth projective variety of dimension $n\geq 1$, and let $H$ be an ample and globally generated line bundle on $X$. A vector bundle $\mathscr{E}$ on $X$ is said to be Ulrich with respect to $H$ if
$H^{i}(X,\mathscr{E}(-jH))=0,$
for all $i\geq 0$ and $1\leq j\leq n$.
Some considerations justify this more general framework. Firstly, numerous essential properties of classical Ulrich bundles remain valid under these assumptions. Secondly, bidouble covers of $\mathbb{P}^{2}$ naturally possess polarizations that are ample and globally generated, but not necessarily very ample. For these reasons and following the convention in [16, Section 1], we call a pair $(X,H)$ a polarized variety if $X$ is a smooth variety and $H$ is an ample and globally generated line bundle.
Definition 2.2.
The smallest $r\geq 1$ such that there is a rank $r$ Ulrich bundle on the polarized variety $(X,H)$ is called the Ulrich complexity of $X$, denoted $\operatorname{uc}(X,H)$.
When focusing on surfaces, the Ulrich condition of a rank $r$ vector bundle $\mathscr{E}$ reduces to two cohomological vanishing together with two conditions on the Chern classes of $\mathscr{E}$, as follows.
Proposition 2.3 ([9, Proposition 2.1]).
Let $(S,H)$ be a smooth polarized surface. For a vector bundle $\mathscr{E}$ on $S$, the following assertions are equivalent:
  1. (1)
    $\mathscr{E}$ is an Ulrich bundle,
  2. (2)
    $\mathscr{E}^{\vee}(3H+K_{S})$ is an Ulrich bundle,
  3. (3)
    $\mathscr{E}$ is an aCM bundle and
    $c_{1}(\mathscr{E})\cdot H$ $=\tfrac{\operatorname{rk}(\mathscr{E})}{2}(3H+K_{S})\cdot H,$ (1)
    $c_{2}(\mathscr{E})$ $=\tfrac{1}{2}\big(c_{1}^{2}(\mathscr{E})-c_{1}(\mathscr{E})\cdot K_{S}\big)-\operatorname{rk}(\mathscr{E})(H^{2}-\chi(\mathscr{O}_{S})),$ (2)
  4. (4)
    $h^{0}(S,\mathscr{E}(-H))=h^{0}(S,\mathscr{E}^{\vee}(2H+K_{S}))=0$ and Equalities (1) and (2) hold.
Definition 2.4.
Let $X$ be a smooth projective variety of dimension $d$, and let $H$ be an ample and globally generated divisor on $X$. A rank $2$ vector bundle $\mathscr{E}$ on $X$ is called special Ulrich if it is Ulrich with respect to $H$ and $c_{1}(\mathscr{E})=(d+1)H+K_{X}$.
Corollary 2.5 ([12, Proposition 5.1.3]).
Let $(S,H)$ be a smooth projective polarized surface and let $\mathscr{E}$ be a rank $2$ vector bundle on $S$. Then $\mathscr{E}$ is a special Ulrich bundle if and only if it satisfies
  1. (1)
    $c_{1}(\mathscr{E})\equiv 3H+K_{S},$
  2. (2)
    $c_{2}(\mathscr{E})=\frac{1}{2}(5H^{2}+3H\cdot K_{S})+2\chi(\mathscr{O}_{S}),$
  3. (3)
    $h^{0}(S,\mathscr{E})>0,$ and
  4. (4)
    $h^{0}(S,\mathscr{E}(-H))=0.$
A vector bundle is called initialized with respect to $H$ if it satisfies the conditions (3) and (4).
A fruitful approach for constructing rank 2 vector bundles on surfaces with prescribed Chern classes is provided by Cayley–Bacharach–type results. For this reason, we state the following theorem, which will play a central role in the proof of Theorem 1.2.
Theorem 2.6 ([26, Theorem 7]).
Let $X$ be a smooth, projective variety of dimension at least $2$, and let $Z\subset X$ be a subscheme of pure codimension $2$. The following statements are equivalent:
  1. (1)
    $Z$ is the zero subscheme of a section of a rank $2$ vector bundle $\mathscr{E}$,
  2. (2)
    there exist three hypersurfaces $E_{1},\,E_{2}$ and $E_{3}$ such that $E_{1}$ and $E_{2}$ have no common components, $Z=E_{1}\cap E_{2}\setminus(E_{1}\cap E_{2}\cap E_{3})$, and $E_{1}\cap E_{2}\cap E_{3}$ is a pure codimension $2$ and Cohen–Maculay.
If (1) or (2) holds, then $\mathscr{E}$ fits into a short exact sequence
$0\longrightarrow\mathscr{O}_{X}\longrightarrow\mathscr{E}\longrightarrow\mathscr{I}_{Z}(E_{1}+E_{2}-E_{3})\longrightarrow 0,$ (3)
which implies $c_{1}(\mathscr{E})=E_{1}+E_{2}-E_{3}$ and $c_{2}({\mathscr{E}})=\#Z$.
Remark 2.7.
The sequence (3) can be found at the end of the proof of [26, Theorem 7].
2.2. Bidouble covers of surfaces
We provide only a brief exposition on bidouble covers on surfaces. For a comprehensive treatment, we refer the reader to [11, 21].
Let $\pi:X\to Y$ be a finite cover of $Y$ with Galois group $G\coloneqq\operatorname{Gal}(\mathbb{C}(X)/\mathbb{C}(Y))\cong(\mathbb{Z}/2\mathbb{Z})^{2}$. Such a map is called a bidouble cover of $Y$. If $X$ is normal, then
$\pi_{*}\mathscr{O}_{X}\cong\mathscr{O}_{Y}\oplus\bigoplus_{i=1}^{3}\mathscr{L}_{\sigma_{i}},$
where $\sigma_{1}$ and $\sigma_{2}$ are generators of $G$ and $\sigma_{3}=\sigma_{1}\sigma_{2}$, and each $\mathscr{L}_{\sigma_{i}}$ is a line bundle on $Y$ determined by $\sigma_{i}$. Thus, $X=\operatorname{Spec}_{\mathscr{O}_{Y}}\left(\pi_{*}\mathscr{O}_{X}\right)$, where the structure of $\mathscr{O}_{S}$-algebra is given by maps $\mathscr{L}_{\sigma_{i}}\oplus\mathscr{L}_{\sigma_{j}}\to\mathscr{L}_{\sigma_{i}+\sigma_{j}}$.
Proposition 2.8 ([11, Proposition 2.3]).
Let $\pi\colon X\to Y$ be a smooth bidouble cover. Then $\pi$ is uniquely determined by six effective divisors $D_{1},D_{2},D_{3}$, and $L_{1},L_{2},L_{3}$ on $Y$ such that
  1. (1)
    $2L_{i}\equiv D_{j}+D_{k}$, and
  2. (2)
    $L_{i}+D_{i}\equiv L_{j}+L_{k}$,
for every permutation $(i,j,k)$ of $(1,2,3)$.
Moreover, the branch locus of $\pi$ is $D=\bigcup_{i=1}^{3}D_{i}$, which must be a normal crossing divisor on $Y$.
Definition 2.9.
A smooth bidouble plane is a smooth bidouble cover $\pi\colon X\to{\mathbb{P}}^{2}$, ramified along three smooth curves $D_{1},D_{2},D_{3}$ of degrees $n_{1},n_{2},n_{3}$. By Proposition 2.8, the relations between the divisors $D_{i}$ and the line bundles $L_{i}$ imply that the integers $n_{1},n_{2},n_{3}$ have the same parity.
We say that $X$ is even (respectively, odd) if each $n_{i}$ is even (respectively, $n_{i}$ is odd).
The following proposition describes several invariants associated with a bidouble plane. A detailed proof, as well as a more general formulation, can be found in [11].
Proposition 2.10.
Let $\pi:X\to\mathbb{P}^{2}$ be a smooth bidouble plane defined as in Proposition 2.8 with $\operatorname{deg}(D_{i})=n_{i}$. Then
  1. (1)
    $\pi_{*}\mathscr{O}_{S}\cong\mathscr{O}_{\mathbb{P}^{2}}\oplus\mathscr{O}_{\mathbb{P}^{2}}{(\frac{-n_{2}-n_{3}}{2})}\oplus\mathscr{O}_{\mathbb{P}^{2}}{(\frac{-n_{1}-n_{3}}{2})}\oplus\mathscr{O}_{\mathbb{P}^{2}}{(\frac{-n_{1}-n_{2}}{2})}$,
  2. (2)
    $K_{S}\equiv\pi^{*}(\mathscr{O}_{\mathbb{P}^{2}}(-3))+R$, where $2R=\pi^{*}(\mathscr{O}_{\mathbb{P}^{2}}(n_{1}+n_{2}+n_{3}))$, and $K_{S}^{2}=(n_{1}+n_{2}+n_{3}-6)^{2}$,
  3. (3)
    $q(S)=0$,
  4. (4)
    $\chi({\mathscr{O}}_{S})=4+\frac{1}{4}(\sum_{i=1}^{3}n^{2}_{i}+n_{1}n_{2}+n_{1}n_{3}+n_{2}n_{3}-6\sum_{i=1}^{3}n_{i})$.
Remark 2.11.
[17, Lemma 2.2] shows that, for a smooth bidouble plane $\pi\colon X\to\mathbb{P}^{2}$, at most one of the branch divisors $D_{i}$ can be trivial; otherwise the surface $S$ would be disconnected.
3. Nerón–Severi group of general bidouble planes
This section aims to compute the Picard number of a bidouble cover. A central feature of the construction of Catanese [11, Section 2] is the existence of three intermediate quotients $Y_{1},Y_{2},Y_{3}$, corresponding to the quotients by the three index-two subgroups of $G$. Concretely, for a bidouble cover $\pi\colon S\to Y$ one has the commutative diagram
[…]
where all the arrows are double covers. The intermediate quotient $Y_{i}$ is branched along $D_{j}+D_{k}$ whenever $(i,j,k)$ is a permutation of $(1,2,3)$.
The lemma below relates the Picard number of $S$ to the Picard numbers of the intermediate quotients. Its proof closely follows [17, Corollary 3.17], which we restate here for clarity.
Lemma 3.1.
Let $\pi\colon S\to{\mathbb{P}}^{2}$ be a smooth bidouble plane. Suppose that $\rho(Y_{2})=\rho(Y_{3})=1$, then $\rho(S)=\rho(Y_{1})$.
Proof.
Let $\sigma_{1}$ and $\sigma_{2}$ be generators of $G$ and let $\sigma_{3}=\sigma_{1}\sigma_{2}$. Let $Y_{i}$ denote the quotient surface $S/\langle\sigma_{i}\rangle$. We can decompose $H^{2}(S,{\mathbb{Q}})$ into four eigenspaces corresponding to the characters of $G$:
$H^{2}(S,\mathbb{Q})=H^{2}(S,{\mathbb{Q}})^{++}\oplus H^{2}(S,{\mathbb{Q}})^{+-}\oplus H^{2}(S,{\mathbb{Q}})^{-+}\oplus H^{2}(S,{\mathbb{Q}})^{--}.$
Here, $H^{2}(S,{\mathbb{Q}})^{\varepsilon_{1}\varepsilon_{2}}$ is the subspace where $\sigma_{i}$ acts by $\varepsilon_{i}\in\{\pm 1\}$. Notice that $H^{2}({\mathbb{P}}^{2},{\mathbb{Q}})\cong H^{2}(S,{\mathbb{Q}})^{++}$ and due to the fact that $H^{2}(S/\langle\sigma_{i}\rangle,{\mathbb{Q}})\cong H^{2}(S,{\mathbb{Q}})^{\langle\sigma_{i}\rangle}$, we have the following isomorphisms:
$H^{2}(Y_{1},\mathbb{Q})$ $\cong H^{2}(S,{\mathbb{Q}})^{++}\oplus H^{2}(S,{\mathbb{Q}})^{+-},$
$H^{2}(Y_{2},\mathbb{Q})$ $\cong H^{2}(S,{\mathbb{Q}})^{++}\oplus H^{2}(S,{\mathbb{Q}})^{-+},$ (5)
$H^{2}(Y_{3},\mathbb{Q})$ $\cong H^{2}(S,{\mathbb{Q}})^{++}\oplus H^{2}(S,{\mathbb{Q}})^{--}.$
Then, we have
$H^{2}(S,\mathbb{Q})=H^{2}(Y_{1},{\mathbb{Q}})\oplus H^{2}(S,{\mathbb{Q}})^{-+}\oplus H^{2}(S,{\mathbb{Q}})^{--}.$
Now, we pass to algebraic classes. Since $S$ is smooth projective, the Néron–Severi group $\operatorname{NS}(S)$ is a Hodge substructure of $H^{2}(S,{\mathbb{Q}})$ stable under $G$. Thus, it decomposes accordingly:
$\operatorname{NS}(S)\otimes{\mathbb{Q}}=(\operatorname{NS}(S)^{++}\oplus\operatorname{NS}(S)^{+-}\oplus\operatorname{NS}(S)^{-+}\oplus\operatorname{NS}(S)^{--})\otimes{\mathbb{Q}}.$
For a finite double cover $p_{i}:S\to Y_{i}$, the pullback induces an injection $p_{i}^{*}:\operatorname{NS}(Y_{i})\otimes{\mathbb{Q}}\hookrightarrow\operatorname{NS}(S)\otimes{\mathbb{Q}}$ with image exactly the $\sigma_{i}$-invariant part $\operatorname{NS}(S)^{\sigma_{i}=+1}$; conversely, any divisor class on $S$ fixed by $\sigma_{i}$ descends to $Y_{i}$. Therefore
$\operatorname{NS}(Y_{1})\otimes{\mathbb{Q}}$ $\cong(\operatorname{NS}(S)^{++}\oplus\operatorname{NS}(S)^{+-})\otimes{\mathbb{Q}},$
$\operatorname{NS}(Y_{2})\otimes{\mathbb{Q}}$ $\cong(\operatorname{NS}(S)^{++}\oplus\operatorname{NS}(S)^{-+})\otimes{\mathbb{Q}},$ (6)
$\operatorname{NS}(Y_{3})\otimes{\mathbb{Q}}$ $\cong(\operatorname{NS}(S)^{++}\oplus\operatorname{NS}(S)^{--})\otimes{\mathbb{Q}}.$
Since $\operatorname{NS}({\mathbb{P}}^{2})={\mathbb{Z}}\cdot H$, $\operatorname{NS}(S)^{++}\cong\pi^{*}\operatorname{NS}({\mathbb{P}}^{2})$ has rank 11. By hypothesis $\rho(Y_{2})=1$ and $\rho(Y_{3})=1$. Comparing ranks in (3) gives
$\rho(Y_{2})=\operatorname{rank}\operatorname{NS}(S)^{++}+\operatorname{rank}\operatorname{NS}(S)^{-+}=1,$
which implies $\operatorname{rank}\operatorname{NS}(S)^{-+}=0$ and analogously one can prove that $\operatorname{rank}\operatorname{NS}(S)^{--}=0$. Hence
$\operatorname{NS}(S)\otimes{\mathbb{Q}}=(\operatorname{NS}(S)^{++}\oplus\operatorname{NS}(S)^{+-})\otimes{\mathbb{Q}}.$
Using (3) we obtain
$\operatorname{NS}(Y_{1})\otimes{\mathbb{Q}}\cong(\operatorname{NS}(S)^{++}\oplus\operatorname{NS}(S)^{+-})\otimes{\mathbb{Q}}\cong\operatorname{NS}(S)\otimes{\mathbb{Q}},$
so $\rho(Y_{1})=\rho(S)$, which gives the desired result. ∎
Remark 3.2.
If $S$ and $Y_{1}$ are smooth and $\rho(Y_{1})=\rho(S)$, then for any line bundle $\mathscr{L}$ on $S$ we have $\pi^{*}(\operatorname{Norm}(\mathscr{L}))\equiv 2\mathscr{L}$. In particular, every element of $\operatorname{NS}(S)/\pi^{*}\operatorname{NS}(Y_{1})$ has order $\leq 2$. The line bundle $\operatorname{Norm}(\mathscr{L})$ on $Y_{1}$, called the norm of $\mathscr{L}$, is defined as follows: for each point $y\in Y_{1}$, the fiber $\operatorname{Norm}(\mathscr{L})_{y}$ is the tensor product $\bigotimes_{x\in\pi^{-1}(y)}\mathscr{L}_{x}^{\otimes n_{\pi,x}}$, where $n_{\pi,x}$ denotes the multiplicity of $\pi$ at the point $x$.
The corollary below follows directly from the proof of Lemma 3.1.
Corollary 3.3.
Let $\pi:S\to{\mathbb{P}}^{2}$ be a smooth bidouble plane. Then $\rho(S)=1$ if and only if
$\rho(Y_{1})=\rho(Y_{2})=\rho(Y_{3})=1.$
The following lemma and proposition provide sufficient conditions for the intermediate quotients to have minimal Picard number.
Lemma 3.4.
Let $\pi\colon Y\to{\mathbb{P}}^{2}$ be the double cover branched along $D=D_{1}+D_{2}$, where $D_{1}$ and $D_{2}$ are general smooth plane curves of degrees $n_{1}$ and $n_{2}$, respectively, meeting transversely, and suppose that $n_{1}+n_{2}\geq 6$. Let $\nu:\widetilde{Y}\to Y$ be the minimal resolution. Then $\rho(\widetilde{Y})=1+n_{1}n_{2}$, and hence $\rho(Y)=1$.
Proof.
Let $H$ be the class of a line in $\mathbb{P}^{2}$, and write $2d=n_{1}+n_{2}\geq 6$. Let $f_{i}\in H^{0}(\mathbb{P}^{2},\mathscr{O}_{\mathbb{P}^{2}}(n_{i}))$ be homogeneous equations defining $D_{i}$, and set
$Y\;=\;\{\,w^{2}=f_{1}(x,y,z)f_{2}(x,y,z)\,\}\subset\mathbb{P}(1,1,1,d).$
For a general choice of $(f_{1},f_{2})$, the curves $D_{1}$ and $D_{2}$ meet transversely in $n_{1}n_{2}$ points and have no other singularities. Consequently, $Y$ has exactly $n_{1}n_{2}$ rational double points (of type $A_{1}$), one over each point $p\in D_{1}\cap D_{2}$. Denote by $E_{p}$ the exceptional (2)(-2)–curve above $p$ in the minimal resolution $\nu:\widetilde{Y}\to Y$.
The classes $(\pi\circ\nu)^{*}H$ and the $E_{p}$ are algebraic on $\widetilde{Y}$. Let $L$ be the lattice generated by $(\pi\circ\nu)^{*}H$ and the $E_{p}$ for every $p\in D_{1}\cap D_{2}$. These classes are linearly independent, thus $\rho(\widetilde{Y})\geq 1+n_{1}n_{2}$. To compute the Picard numbers of $\widetilde{Y}$ and $Y$, consider a one-parameter family $f_{t}\in H^{0}(\mathbb{P}^{2},\mathscr{O}_{\mathbb{P}^{2}}(2d))$ with $f_{0}=f_{1}f_{2}$ and such that for $t\neq 0$, the surface $Y_{t}=\{w^{2}=f_{t}\}$ is smooth. Define
$\mathscr{Y}=\{\,w^{2}=f_{t}(x,y,z)\,\}\;\subset\;\mathbb{P}(1,1,1,d)\times\Delta,$
where $\Delta$ is a small analytic disk around $0\in\mathbb{C}$. Let $\pi_{\Delta}:\mathscr{Y}\to\Delta$ be the projection to the parameter. Then $\mathscr{Y}\to\Delta$ is flat, with central fiber $Y_{0}=Y$ and smooth fibers $\pi_{t}\colon Y_{t}\to{\mathbb{P}}^{2}$ for $t\neq 0$.
By the semistable reduction theorem (see [20, p. 102]), there exists a finite base change $\Delta^{\prime}\to\Delta$ (ramified only over 0) and a birational modification of the total space such that the induced family $\mathscr{Y}^{\prime}\to\Delta^{\prime}$ is semistable, that is, its central fiber is a normal crossings divisor with smooth components. Replacing $\mathscr{Y}$ by this semistable model ensures that the Clemens–Schmid sequence and the specialization morphism of mixed Hodge structures apply functorially. In the semistable model, besides the main component $\widetilde{Y}$, the central fiber contains, for each node $p\in D_{1}\cap D_{2}$, an additional smooth rational surface $F_{p}$, obtained as the minimal resolution of the corresponding local double point degeneration. Moreover, each $F_{p}$ meets $\widetilde{Y}$ transversely along the exceptional curve $E_{p}$.
By Buium [6, Theorem, p. 1], a general double plane branched along a smooth divisor of degree at least six has Picard number $1$. Hence, the orthogonal complement (with respect to the cup product) of $\pi_{t}^{*}H$ in $H^{2}(Y_{t},\mathbb{Q})$, denoted $P_{t}$, satisfies $P_{t}\cap H^{1,1}(Y_{t})=0$ for $t$ sufficiently close to 0.
[…]
The kernel of $\mathrm{sp}$ is generated by the vanishing cycles, which on $t=0$ correspond to the classes of the exceptional curves $E_{p}$ by the Clemens-Schmid exact sequence (see [20, p. 108]). Since $\alpha\in P$ is orthogonal to each $E_{p}$, we have $\mathrm{sp}(\alpha)\neq 0$.
The hyperplane class $\pi_{t}^{*}H$ is the restriction of ${\mathscr{O}}_{{\mathbb{P}}(1,1,1,d)}(1)$ to $Y_{t}$. Therefore,
$\langle\mathrm{sp}(\alpha),\pi_{t}^{*}H\rangle=\langle\alpha,(\pi\circ\nu)^{*}H\rangle=0.$
Thus $\mathrm{sp}(\alpha)\in P_{t}\cap H^{1,1}(Y_{t})$, which is a contradiction as $\mathrm{sp}(\alpha)\neq 0$. This contradiction proves that $P\cap H^{1,1}(\widetilde{Y})=0$ as claimed. Therefore $\mathrm{NS}(\widetilde{Y})\otimes\mathbb{Q}=L_{\mathbb{Q}}$, which implies that $\rho(\widetilde{Y})=1+n_{1}n_{2}$. Finally, since $Y$ has only rational double points, the natural map $\mathrm{Pic}(\widetilde{Y})\to\mathrm{Pic}(Y)$ is surjective with kernel generated by the exceptional curves $\{E_{p}\}$. Consequently $\rho(Y)=\rho(\widetilde{Y})-n_{1}n_{2}=1$, which yields the desired result. ∎
Proposition 3.5.
Let $Y$ be a double cover of ${\mathbb{P}}^{2}$ branched along $D=D_{1}+D_{2}$ where $D_{1}$ and $D_{2}$ are general effective divisors of degrees $n_{1}$ and $n_{2}$, respectively, and assume that $n_{1}\leq n_{2}$. Then $\rho(Y)>1$ if and only if $(n_{1},n_{2})\in\{(0,2),(0,4),(1,3),(2,2)\}$.
Proof.
By Lemma 3.4, we have that $\rho(Y)=1$ whenever $n_{1}+n_{2}\geq 6$.
Assume $n_{1}+n_{2}=4$. The resolution $\widetilde{Y}$ of $Y$ is a rational surface with Picard number $\rho(\widetilde{Y})=\chi({\mathscr{O}}_{\widetilde{Y}})-2=8$. Since $\widetilde{Y}$ is obtained by blowing up $Y$ at $n_{1}n_{2}$ points, we have $\rho(\widetilde{Y})=\rho(Y)+n_{1}n_{2}$. Hence $\rho(Y)=8$ for $(n_{1},n_{2})=(0,4)$, $\rho(Y)=5$ for $(n_{1},n_{2})=(1,3)$, and $\rho(Y)=4$ for $(n_{1},n_{2})=(2,2)$.
Finally, suppose that $n_{1}+n_{2}=2$. When $(n_{1},n_{2})=(0,2)$, $Y$ is isomorphic to ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$, then $\rho(Y)=2$. When $(n_{1},n_{2})=(1,1)$, $Y$ is a quadratic cone in ${\mathbb{P}}^{3}$, then $\rho(Y)=1$. ∎
Proof of Theorem 1.1.
By Corollary 3.3, $S$ has Picard number bigger than 11 if and only if some intermediate quotient has Picard number bigger than 11. Proposition 3.5 implies that some intermediate quotient has Picard number bigger than 11 if and only if some permutation of $(n_{1},n_{2},n_{3})$ belongs to
$\{(0,2,2n)\colon n\geq 1\}\cup\{(0,4,2n)\colon n\geq 2\}\cup\{(1,3,2n+1)\colon n\geq 0\}\cup\{(2,2,2n)\colon n\geq 1\}.$
This yields the desired result. ∎
4. Ulrich line bundles on general bidouble Planes
In this section, we determine the range of branch degrees for which Ulrich line bundles could exist. The strategy is to analyze the existence of Ulrich line bundles on smooth bidouble planes by distinguishing two complementary cases, as dictated by Theorem 1.1. First, we establish non-existence results for covers branched in sufficiently high degree, and then we turn to the low-degree cases where existence may occur. Throughout this section, we denote by $H$ the line bundle $\pi^{*}\mathscr{O}_{\mathbb{P}^{2}}(1)$.
4.1. Non-existence of Ulrich line bundle in high degree
We begin by showing that, for bidouble covers of ${\mathbb{P}}^{2}$ branched along divisors of large degree, Ulrich line bundles cannot exist.
Lemma 4.1.
Let $\pi\colon S\to{\mathbb{P}}^{2}$ be a smooth odd bidouble plane. Then, $S$ admits no Ulrich vector bundles of odd rank with respect to $H$.
Proof.
Let $n_{1}$, $n_{2}$, and $n_{3}$ be the ramification degrees of $\pi$, and let
$n=n_{1}+n_{2}+n_{3}.$
Since each $n_{i}$ is odd, the sum $n$ is also odd. By Proposition 2.10, we have $2K_{S}=(n-6)H$. For the sake of contradiction, assume that an odd rank Ulrich bundle $\mathscr{L}$ on $S$ exists with respect to $H$. By Proposition 2.3, in particular, we have
$c_{1}(\mathscr{L})\cdot H\;=\;\frac{\operatorname{rk}({\mathscr{L}})}{2}(3H+K_{S})\cdot H.$
Using $H^{2}=4$ and $K_{S}\cdot H=2(n-6)$, and multiplying both sides by n6n-6, we obtain
$2c_{1}(\mathscr{L})\cdot K_{S}$ $=\frac{\operatorname{rk}({\mathscr{L}})(n-6)}{2}(3H^{2}+K_{S}\cdot H)$
$=n(n-6)\operatorname{rk}({\mathscr{L}}).$
Since the left-hand side is an integer, the right-hand side must also be an integer. But $n(n-6)\operatorname{rk}({\mathscr{L}})$ is odd, while $2\,c_{1}(\mathscr{L})\cdot K_{S}$ is even, which is a contradiction. Thus, no odd rank Ulrich bundle can exist. ∎
Lemma 4.2.
Let $\pi\colon S\to{\mathbb{P}}^{2}$ be a smooth even bidouble plane that admits an Ulrich line bundle with respect to $H$. Then the Picard number $\rho(S)$ satisfies $\rho(S)>1$.
Proof.
Let $n_{1}$, $n_{2}$, and $n_{3}$ be the ramification degrees of $\pi$. Assume to the contrary that $S$ admits an Ulrich line bundle $\mathscr{L}$} and that $\rho(S)=1$. Then $\pi^{*}(\operatorname{Pic}({\mathbb{P}}^{2}))$ is a finite-index subgroup of $\operatorname{Pic}(S)$, so there exist line bundles ${\mathscr{L}}^{\prime}\in\operatorname{Pic}(S)$ and ${\mathscr{T}}\in\operatorname{Pic}(S)_{tor}$ integers $a$ and $m\geq 1$, with $\gcd(a,m)=1$ such that ${\mathscr{L}}\cong{\mathscr{L}}^{\prime}\otimes{\mathscr{T}}$ and $m{\mathscr{L}}^{\prime}={\mathscr{O}}_{S}(a)$. Notice that $\mathscr{L}$} and $\mathscr{L}$$^{\prime}$ are numerically equivalent. Since
$c_{1}({\mathscr{L}})\cdot H=\frac{4a}{m},$
and $\gcd(a,m)=1$, it follows that $m$ divides $4$. Thus, $m\in\{1,2,4\}$. We claim that $m=1$.
From $K_{S}\cdot H=2(n_{1}+n_{2}+n_{3})-12$ and Proposition 2.3, we obtain
$\frac{4a}{m}=c_{1}(\mathscr{L})\cdot H\;=\;\tfrac{1}{2}(3H+K_{S})\cdot H=n_{1}+n_{2}+n_{3}.$ (7)
If $m=4$, then $a=n_{1}+n_{2}+n_{3}$, an even number, contradicting $\gcd(a,m)=1$.
Now assume $m=2$. Proposition 2.3 gives
$0=\frac{1}{2}(c_{1}^{2}({\mathscr{L}})-c_{1}({\mathscr{L}})\cdot K_{S})+\chi(\mathscr{O}_{S})-H^{2}.$ (8)
Hence $c_{1}^{2}({\mathscr{L}})-c_{1}({\mathscr{L}})\cdot K_{S}$ is even.
Since
$c_{1}^{2}({\mathscr{L}})=\frac{4a^{2}}{m^{2}}\qquad\text{and}\qquad c_{1}({\mathscr{L}})\cdot K_{S}=\frac{2a}{m}(n_{1}+n_{2}+n_{3}-6),$ (9)
we obtain
$c_{1}^{2}({\mathscr{L}})-c_{1}({\mathscr{L}})\cdot K_{S}=a^{2}-a(n_{1}+n_{2}+n_{3}-6).$
Because $n_{1}+n_{2}+n_{3}$ is even, the above expression is even only if $a$ is even, again contradicting $\gcd(a,m)=1$. Thus $m\neq 2$, and therefore $m=1$.
We conclude that any Ulrich line bundle $\mathscr{L}$} must be numerically equivalent to some divisor ${\mathscr{O}}_{S}(a)$ for some integer $a$. Moreover, (7) yields $4a=n_{1}+n_{2}+n_{3}$. Using (9) with $m=1$ and the expression for $\chi({\mathscr{O}}_{S})$ given in Proposition 2.10, the Ulrich condition (8) becomes
$0=\frac{1}{2}\left(3(n_{1}+n_{2}+n_{3})-\frac{(n_{1}+n_{2}+n_{3})^{2}}{4}\right)+\frac{1}{4}\left((n^{2}_{1}+n^{2}_{2}+n^{2}_{3})+n_{1}n_{2}+n_{1}n_{3}+n_{2}n_{3}-6(n_{1}+n_{2}+n_{3})\right),$
hence $n_{1}^{2}+n_{2}^{2}+n_{3}^{2}=0$, which is a contradiction. Therefore $\rho(S)\neq 1$, hence $\rho(S)>1$. ∎
Theorem 1.1, Lemma 4.1, and Lemma 4.2 have the following consequence.
Corollary 4.3.
Let $S\to\mathbb{P}^{2}$ be a smooth bidouble plane branched along general curves $D_{1},D_{2},D_{3}$ of degrees $n_{1},n_{2},n_{3}$. If no permutation of $(n_{1},n_{2},n_{3})$ lies in
$\{(0,2,2n):n\geq 1\}\;\cup\;\{(0,4,2n):n\geq 2\}\;\cup\;\{(2,2,2n):n\geq 1\},$
then $S$ admits no Ulrich line bundles with respect to $H$.
4.2. Ulrich line bundles in low degree
We now turn to the complementary situation where at least one pair of branch degrees is small. In this range, the Picard number can jump, and one might hope for the existence of Ulrich line bundles.
For double covers of ${\mathbb{P}}^{2}$, it was shown in [22] that the only obstruction is having Picard number $1$. In contrast, we will exhibit families of bidouble covers with Picard number greater than one that nonetheless admit no Ulrich line bundles. At the same time, we also identify particular low-degree configurations in which Ulrich line bundles exist.
Proposition 4.4.
Let $\pi\colon S\to{\mathbb{P}}^{2}$ be a bidouble plane branched along the union of the zero divisor, a smooth conic, and a smooth curve of degree $2n$ for some $n\geq 3$, all in general position. Then $S$ admits no Ulrich line bundle with respect to $H$.
Proof.
Notice that there is a morphism $\pi_{1}\colon S\to{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$, which is a double cover branched along the divisor $(2n,2n)$. For the sake of contradiction, suppose that there exists an Ulrich line bundle $\mathscr{L}$} with respect to $H$. By Proposition 2.3, we have
$c_{1}(\mathscr{L})\cdot H\;$ $=\;\tfrac{1}{2}(3H+K_{S})\cdot H.$ (10)
$0$ $=\frac{1}{2}(c_{1}^{2}({\mathscr{L}})-c_{1}({\mathscr{L}})\cdot K_{S})+\chi(\mathscr{O}_{S})-H^{2}.$ (11)
By Proposition 2.10, $\chi({\mathscr{O}}_{S})=n^{2}-2n+2$, $K_{S}\cdot H=4(n-2)$, and $H^{2}=4$. The Lemma 3.1 and Proposition 3.5 imply $\rho(S)=\rho({\mathbb{P}}^{1}\times{\mathbb{P}}^{1})$, so Remark 3.2 implies that there exists an integer $m\in\{1,2\}$ such that $m{\mathscr{L}}$ is the pullback of a class in $\operatorname{NS}({\mathbb{P}}^{1}\times{\mathbb{P}}^{1})$. Then, there exist integers $a$ and $b$ with
$m{\mathscr{L}}=\pi_{1}^{*}({\mathscr{O}}_{{\mathbb{P}}^{1}\times{\mathbb{P}}^{1}}(a,b)),$
and consequently, we have that
$c_{1}({\mathscr{L}})\cdot H=\frac{2}{m}(a+b),\quad c_{1}({\mathscr{L}})\cdot K_{S}=\frac{2}{m}(n-2)(a+b),\quad\text{and }\quad c_{1}^{2}({\mathscr{L}})=\frac{4ab}{m^{2}}.$
Equation (10) yields $a+b=(n+1)m$, and (11) gives
$0=\frac{1}{2}\left(\frac{4ab}{m^{2}}-\frac{2(n-2)(a+b)}{m}\right)+(n^{2}-2n+2)-4=\frac{2ab}{m^{2}}-n.$
In particular, $2ab=nm^{2}$. Substituting b=(n+1)mab=(n+1)m-a, we have the following quadratic equation in $a$
$2a^{2}-2m(n+1)a+m^{2}n=0.$
The discriminant of this equation is $4m^{2}(n^{2}+1)$, which is never a perfect square for $n\geq 1$. This contradicts the integrality of $a$. ∎
Theorem 4.5 (Ulrich line bundles).
Let $\pi\colon S\to\mathbb{P}^{2}$ be a bidouble plane branched along general curves $D_{1},D_{2},D_{3}$ of degrees $n_{1},n_{2},n_{3}$. If $S$ admits an Ulrich line bundle with respect to $H$, then some permutation of $(n_{1},n_{2},n_{3})$ belongs to
$\{(0,2,2)\}\;\cup\;\{(0,4,2n):n\geq 1\}\;\cup\;\{(2,2,2n):n\geq 1\}.$
Proof.
Corollary 4.3 shows that if a bidouble cover $S$ branched along curves of degrees $(n_{1},n_{2},n_{3})$ admits an Ulrich line bundle, then some permutation of $(n_{1},n_{2},n_{3})$ belongs to
$\{(0,2,2n)\colon n\geq 1\}\;\cup\;\{(0,4,2n)\colon n\geq 2\}\;\cup\;\{(2,2,2n)\colon n\geq 1\}.$
By Proposition 4.4, if a permutation of $(n_{1},n_{2},n_{3})$ is equal to $(0,2,2n)$ with $n\geq 3$, then $S$ admits no Ulrich line bundles with respect to $H$. Therefore, the only possible cases occur when some permutation of $(n_{1},n_{2},n_{3})$ belongs to
$\{(0,2,2)\}\;\cup\;\{(0,4,2n)\colon n\geq 1\}\;\cup\;\{(2,2,2n)\colon n\geq 1\},$
which is precisely the statement of the theorem. This completes the proof. ∎
While the converse of Theorem 4.5 remains open, the following proposition provides explicit examples of bidouble planes carrying Ulrich line bundles.
Proposition 4.6.
Let $\pi\colon S\to\mathbb{P}^{2}$ be a bidouble plane branched along general curves $D_{1},D_{2},D_{3}$ of degrees $n_{1},n_{2},n_{3}$ with $n_{1}\leq n_{2}\leq n_{3}$. If $(n_{1},n_{2},n_{3})$ is $(0,2,2)$ or $(0,2,4)$, $S$ admits an Ulrich line bundle with respect to $H$.
Proof.
We treat the two cases separately.
Case (0,2,2). In this case, $S$ is a del Pezzo surface of degree $4$, and $K_{S}=H$. By [24, Theorem 1.1], every del Pezzo surface of degree $d$ admits an Ulrich line bundle with respect to $K_{X}$. More precisely, a line bundle $\mathscr{L}$ is Ulrich if and only if $\mathscr{L}\cong\mathscr{O}_{S}(D)$ for some rational normal curve $D\subset S$ of degree at most $d$. Therefore, $S$ admits an Ulrich line bundle with respect to $H$.
Case (0,2,4). Let $\pi\colon S\to{\mathbb{P}}^{2}$ be the bidouble plane ramified along the zero divisor, a smooth conic $C$, and a smooth quartic $Q$. Then $S$ is a K3 surface, and following the notation in the diagram (4), it admits two intermediate double planes: $Y_{3}$, the double plane branched along $C$, and $Y_{2}$, the double plane branched along $Q$. The surface $Y_{3}$ is isomorphic to ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$, hence $\rho(Y_{3})=2$. The surface $Y_{2}$ is a del Pezzo surface of degree $2$, so $\rho(Y_{2})=8$, and we may view $Y_{2}$ as the blow–up of ${\mathbb{P}}^{2}$ at $7$ general points. In particular,
$\operatorname{NS}(Y_{2})\supset\langle H_{Y_{2}},E_{1},\dots,E_{7}\rangle,$
where $H_{Y_{2}}$ is the pullback of a line in ${\mathbb{P}}^{2}$ via $\pi_{2}$, and $E_{i}$ are the exceptional divisors.
By Proposition 2.3, a line bundle $\mathscr{L}$} on $S$ is Ulrich with respect to $H$ if and only if
$c_{1}^{2}({\mathscr{L}})=4,\qquad c_{1}({\mathscr{L}})\cdot H=6,\qquad h^{0}(S,{\mathscr{L}}(-H))=h^{0}(S,{\mathscr{L}}^{\vee}(2H))=0.$ (12)
We now construct such a line bundle. Let $\ell\subset{\mathbb{P}}^{2}$ be a line tangent to $C$ which intersects $Q$ in $4$ distinct points. By [22, Theorem 4.4], the preimage $\pi^{-1}(\ell)=\Gamma_{1}+\Gamma_{2}$ is the union of two smooth curves of genus $1$. In particular, $\Gamma_{i}^{2}=2g(\Gamma_{i})-2=0$. Moreover, $\pi^{*}\ell\sim H$, hence $H=\Gamma_{1}+\Gamma_{2}$, and by symmetry, we obtain $H\cdot\Gamma_{j}=2$ for each $j$.
Let $E_{i}^{\prime}:=p_{2}^{*}E_{i}$ denote the pullback of the exceptional divisors from $Y_{2}$ to $S$. Since $H_{Y_{2}}=-K_{Y_{2}}$, we have $H=p_{2}^{*}(H_{Y_{2}})=p_{2}^{*}(-K_{Y_{2}})$. In particular,
$H\cdot E_{i}^{\prime}=p_{2}^{*}(-K_{Y_{2}})\cdot p_{2}^{*}E_{i}=2(-K_{Y_{2}}\cdot E_{i})=2$
for all $i$. Since $H=\Gamma_{1}+\Gamma_{2}$ we have
$\Gamma_{1}\cdot E_{i}^{\prime}=\Gamma_{2}\cdot E_{i}^{\prime}=1$
for all $i$.
Choose two different exceptional divisors $E_{1},E_{2}$ on $Y_{2}$ and consider the divisor
$D:=H+\Gamma_{1}+E_{1}^{\prime}-E_{2}^{\prime}\in\operatorname{Div}(S).$
First, we verify that the numerical conditions in (12) are satisfied. Using the intersection numbers above and $(E_{i}^{\prime})^{2}=2E_{i}^{2}=-2$, we compute
$D\cdot H=(H+\Gamma_{1}+E_{1}^{\prime}-E_{2}^{\prime})\cdot H=H^{2}+\Gamma_{1}\cdot H+E_{1}^{\prime}\cdot H-E_{2}^{\prime}\cdot H=4+2+2-2=6,$
and
$D^{2}$ $=H^{2}+\Gamma_{1}^{2}+(E_{1}^{\prime})^{2}+(E_{2}^{\prime})^{2}+2\big(H\cdot\Gamma_{1}+H\cdot E_{1}^{\prime}-H\cdot E_{2}^{\prime}+\Gamma_{1}\cdot E_{1}^{\prime}-\Gamma_{1}\cdot E_{2}^{\prime}-E_{1}^{\prime}\cdot E_{2}^{\prime}\big)$
$=4+0-2-2+2\big(2+2-2+1-1-0\big)=4.$
Thus $D^{2}=4$ and $D\cdot H=6$.
We now show that the cohomological vanishings in (12) hold. Due to $H=\Gamma_{1}+\Gamma_{2}$, we have
$D-H=\Gamma_{1}+E_{1}^{\prime}-E_{2}^{\prime},\qquad 2H-D=\Gamma_{2}-E_{1}^{\prime}+E_{2}^{\prime}.$
For the sake of clarity, we show that $h^{0}(S,D-H)=0$; the vanishing $h^{0}(S,2H-D)=0$ is proved similarly. For a contradiction, assume that $h^{0}(S,D-H)>0$, therefore, there exists an effective divisor $F\sim D-H$. A direct computation gives $F^{2}=-4$ and $H\cdot F=2$. Moreover,
$F\cdot E_{1}^{\prime}=\Gamma_{1}\cdot E_{1}^{\prime}+(E_{1}^{\prime})^{2}-E_{2}^{\prime}\cdot E_{1}^{\prime}=1-2-0=-1.$
Since $F$ is effective, the inequality $F\cdot E_{1}^{\prime}<0$ forces $E_{1}^{\prime}$ to be a component of $F$, so we could write $F=E_{1}^{\prime}+F^{\prime}$, with $F^{\prime}$ effective. However, notice that $(F^{\prime})^{2}=(\Gamma_{1}-E_{2}^{\prime})^{2}=-4$ and $H\cdot F^{\prime}=H\cdot\Gamma_{1}-H\cdot E_{2}^{\prime}=0$. Since $H$ is ample, any nonzero effective divisor $Z$ must satisfy $H\cdot Z>0$, hence $F^{\prime}$ cannot be effective. This contradiction shows $h^{0}(S,D-H)=0$.
Putting everything together, the line bundle ${\mathscr{L}}:={\mathscr{O}}_{S}(D)$ satisfies
$c_{1}^{2}({\mathscr{L}})=4,\qquad H\cdot c_{1}({\mathscr{L}})=6,\qquad h^{0}(S,{\mathscr{L}}(-H))=h^{0}(S,{\mathscr{L}}^{\vee}(2H))=0,$
so, by Proposition 2.3, $\mathscr{L}$} is an Ulrich line bundle with respect to $H$. ∎
Remark 4.7.
In case $(2,2,2)$, [17, Theorem 4.8] implies that $S$ is an Enriques surface of degree $4$. In this case, $S$ is nodal, and therefore the methods of [5] do not apply.
In case $(0,4,4)$, one can construct a divisor analogous to the one used in case $(0,2,4)$, which satisfies the numerical conditions of Proposition 2.3. However, verifying the required cohomological vanishings appears to be substantially more delicate, and we were unable to establish them. More precisely, instead of choosing a line tangent to the conic as in the case $(0,2,4)$, one may consider a smooth conic whose intersection with one of the quartic branch curves is an even divisor. The existence of such a conic follows from [22, Corollary 1.3]. Nevertheless, this approach does not seem to lead to a proof of the necessary vanishing conditions.
For the remaining cases of Theorem 4.5 that could a priori admit an Ulrich line bundle, Lemma 3.1 shows that $\rho(S)=\rho(Y_{3})$. This strongly restricts the geometry of $\operatorname{Pic}(S)$: in particular, for any line bundle $\mathscr{L}\in\operatorname{Pic}(S)$, a suitable multiple of $\mathscr{L}$ is the pullback of a line bundle on $Y_{3}$. This observation suggests that the existence of Ulrich line bundles on $S$ in these cases is unlikely, as illustrated by Proposition 4.4.
5. rank two Ulrich bundles on general even bidouble planes
In this section, we prove Theorem 1.2. We show that, in contrast to the situation for Ulrich line bundles, special Ulrich bundles exist for almost every even generic bidouble cover $S$ of the projective plane. Recall that $H$ denotes the polarization $\pi^{*}\mathscr{O}_{{\mathbb{P}}^{2}}(1)$.
By Corollary 2.5, in order to construct a special Ulrich bundle with respect to $H$, it suffices to produce a rank-two vector bundle $\mathscr{E}$ satisfying
$c_{1}(\mathscr{E})\equiv 3H+K_{S},\qquad c_{2}(\mathscr{E})=\frac{1}{2}(5H^{2}+3H\cdot K_{S})+2\chi(\mathscr{O}_{S}),$
together with the cohomological conditions
$h^{0}(S,\mathscr{E})>0,\qquad h^{0}(S,\mathscr{E}(-H))=0.$
The key idea is to apply Theorem 2.6 to construct a vector bundle of rank two satisfying these requirements. More precisely, we choose hypersurfaces $E_{1}$, $E_{2}$, $E_{3}$ such that
$3H+K_{S}\equiv E_{1}+E_{2}-E_{3},\qquad\text{and}\qquad\#Z=\frac{1}{2}(5H^{2}+3H\cdot K_{S})+2\chi(\mathscr{O}_{S}),$
where $Z=(E_{1}\cap E_{2})\setminus(E_{1}\cap E_{2}\cap E_{3})$. According to Theorem 2.6, these data produce a rank-two vector bundle $\mathscr{E}$ with the desired Chern classes. The remainder of the proof consists in using the short exact sequence (3) to verify the required cohomology vanishings.
Theorem 5.1 (Ulrich bundles of rank two).
Let $\pi\colon S\to\mathbb{P}^{2}$ be an even smooth bidouble plane branched along smooth curves $D_{1},D_{2},D_{3}$ of degrees $n_{1},n_{2},n_{3}$, respectively. If no permutation of $(n_{1},n_{2},n_{3})$ is equal to $(0,2,2)$, then $S$ carries a special Ulrich bundle $\mathscr{E}$ with respect to $H$.
Proof.
Let $_{i}=n_{i}/2, m=m$$m_{1}+m_{2}+m_{3}$ and $M=m^{2}+m_{1}^{2}+m_{2}^{2}+m_{3}^{2}$, and notice that $M$ is even. By Corollary 2.5, a rank-two vector bundle $\mathscr{E}$ is a special Ulrich bundle with respect to $H$ if and only if it satisfies
$c_{1}(\mathscr{E})\equiv mH,\qquad c_{2}(\mathscr{E})=M,\qquad h^{0}(S,\mathscr{E})>0,\qquad\text{and}\qquad h^{0}(S,\mathscr{E}(-H))=0.$
Now, we define the hypersurfaces $E_{1}$, $E_{2}$, and $E_{3}$ according to the reduction of $M$ modulo $4$.
If $M\equiv 0\pmod{4}$, set $E_{1}\colon=\pi^{*}(\ell)$, $E_{2}\colon=\pi^{*}(C)$, and $E_{3}\colon=\pi^{*}(C^{\prime})$, where $\ell$ is a general line, and $C$ and $C^{\prime}$ are smooth plane curves in general position of degrees $M/4$, and M/4+1mM/4+1-m, respectively.
On the other hand, if $M\equiv 2\pmod{4}$, set $E_{2}\colon=\pi^{*}(C)$, and $E_{3}\colon=\pi^{*}(C^{\prime})$, where $C$ and $C^{\prime}$ are smooth plane curves in general position of degrees $(M+2)/4$, and (M+2)/4+1m(M+2)/4+1-m, respectively, and $E_{1}\colon=\pi^{*}(\ell)$, where $\ell$ is a line that intersects exactly one point of $C$ and the $D_{i}$’s.
In any case, since $m\geq 3$, $\deg(C)\geq m$, in particular, the degree of $C^{\prime}$ is positive. Due to the generality, we have that $E_{1},\,E_{2}$, and $E_{3}$ have no common point.
Let $Z\subset S$ be the zero-dimensional subscheme
$Z\coloneqq E_{1}\cap E_{2}\;\setminus\;(E_{1}\cap E_{2}\cap E_{3}).$
Since the triple intersection is empty and because of the choice of $E_{i}$’s, if $M\equiv 0\pmod{4}$, we have $Z=4\left(\frac{M}{4}\right)=M$. In the case $M\equiv 2\pmod{4}$, there is a single point in $\ell\cap C$ which if ramified via $\pi$, then $Z=4\left(\frac{M+2}{4}-1\right)+2=M$. By construction, the data $(E_{1},E_{2},E_{3},Z)$ satisfy Theorem 2.6 (2). Hence, there exists a rank two vector bundle $\mathscr{E}$ fitting into
$0\longrightarrow\mathscr{O}_{S}\longrightarrow\mathscr{E}\longrightarrow\mathscr{I}_{Z}(E_{1}+E_{2}-E_{3})\longrightarrow 0.$ (13)
Since $E_{1}+E_{2}-E_{3}\equiv mH$, we obtain $c_{1}(\mathscr{E})\equiv mH$, and $c_{2}(\mathscr{E})=\#Z=M$.
Now, we claim that $h^{0}(S,\mathscr{E}(-H))=0$. Twisting (13) by $\mathscr{O}_{S}(-H)$ gives
$0\longrightarrow\mathscr{O}_{S}(-H)\longrightarrow\mathscr{E}(-H)\longrightarrow\mathscr{I}_{Z}((m-1)H)\longrightarrow 0.$ (14)
Since $h^{0}(S,{\mathscr{O}}_{S}(-H))=0$, it is enough to prove that $h^{0}(S,\mathscr{I}_{Z}((m-1)H))=0$.
For the sake of contradiction, assume that there exists a nonzero section of $\mathscr{I}_{Z}((m-1)H)$, and let $D\in|(m-1)H|$ be the corresponding effective divisor. Then $Z\subset D$. Intersecting with $E_{1}\in|H|$, we obtain $D\cdot E_{1}=4(m-1)$. On the other hand, since $Z\subset E_{1}$ and every point of $Z$ contributes at least one to the intersection multiplicity, we have $D\cdot E_{1}\;\geq\;\#Z=M$. Because $m_{1}^{2}+m_{2}^{2}+m_{3}^{2}\geq\frac{1}{3}(m_{1}+m_{2}+m_{3})^{2}=\frac{1}{3}m^{2}$, we obtain $M\geq 4m^{2}/3$. For $m\geq 3$, this gives $M>4(m-1)=D\cdot E_{1}$, a contradiction. Therefore $h^{0}\bigl(S,\mathscr{I}_{Z}((m-1)H)\bigr)=0$, and hence $h^{0}(S,\mathscr{E}(-H))=0$.
Finally, the inclusion $\mathscr{O}_{S}\hookrightarrow\mathscr{E}$ induces a nonzero section $s\in H^{0}(S,\mathscr{E})$. Combining all the above, the vector bundle $\mathscr{E}$ satisfies
$c_{1}(\mathscr{E})\equiv mH,\qquad c_{2}(\mathscr{E})=M,\qquad h^{0}(S,\mathscr{E})>0,\qquad\text{and}\qquad h^{0}(S,\mathscr{E}(-H))=0,$
as desired. ∎
Proof of Theorem 1.2.
Let us recall the definition of $T_{1}$ and $T_{2}$:
$T_{1}=\{(0,4,2n)\mid n\geq 2\}\;\cup\;\{(2,2,2n)\mid n\geq 1\},\qquad\text{and}\qquad T_{2}=\{(0,2,2),\,(0,2,4)\}.$
To begin with, assume that $S$ is an odd bidouble cover. Theorem 4.5 implies that $S$ does not admit an Ulrich line bundle with respect to $H$. Consequently, $\operatorname{uc}(S,H)>1$.
Now, assume that $S$ is an even bidouble cover. First, Suppose that $(n_{1},n_{2},n_{3})\notin T_{1}\cup T_{2}$. Theorem 4.5 implies that $S$ does not admit an Ulrich line bundle with respect to $H$. On the other hand, Theorem 5.1 implies the existence of an Ulrich rank two vector bundle. Consequently, $\operatorname{uc}(S,H)=2$. Now, if $(n_{1},n_{2},n_{3})\in T_{1}$ Theorem 5.1 implies that $\operatorname{uc}(S,H)\leq 2$. Finally, if $(n_{1},n_{2},n_{3})\in T_{2}$ Proposition 4.6 implies that $S$ admits an Ulrich line bundle with respect to $H$, which yields $\operatorname{uc}(S,H)=1$. ∎

### v2

@@
-Ulrich complexity and Picard rank of bidouble Planes
+Picard rank and Ulrich line bundles on bidouble planes
 Abstract.
@@
 For a bidouble plane branched along three smooth curves of degrees $n_{1},n_{2},n_{3}$, we show that $\rho(S)=1$ unless $(n_{1},n_{2},n_{3})$ belongs to an explicit list, thereby extending Buium’s classical results on double planes to the non-cyclic case.
-As an application, we determine the range of branch degrees for which Ulrich line bundles could exist, and we show that every even bidouble plane carries a special rank two Ulrich bundle.
+As an application, we determine the range of branch degrees for which Ulrich line bundles could exist.
 Our method combines the invariant-theoretic decomposition of $H^{2}(S,{\mathbb{Q}})$ under the Galois group with cohomological criteria for Ulrich bundles.
@@
 Theorem 4.5).
-We therefore turn to the existence of rank two Ulrich bundles on $S$.
 The smallest integer $r\geq 1$ such that there exists a rank $r$ Ulrich bundle on $S$ with respect to $H$ is called the Ulrich complexity and is denoted by $\operatorname{uc}(S,H)$.
-Our second main theorem establishes the Ulrich complexity of even bidouble planes.
+Our second main theorem establishes lower bounds on the Ulrich complexity of bidouble planes.
 Theorem 1.2.
@@
   • (a)
-    If $(n_{1},n_{2},n_{3})\notin T_{1}\cup T_{2}$, then $\operatorname{uc}(S,\mathscr{O}_{S}(1))=2$.
+    If $(n_{1},n_{2},n_{3})\notin T_{1}\cup T_{2}$, then $\operatorname{uc}(S,\mathscr{O}_{S}(1))>1$.
   • (b)
-    If $(n_{1},n_{2},n_{3})\in T_{1}$, then $\operatorname{uc}(S,\mathscr{O}_{S}(1))\leq 2$.
-  • (c)
     If $(n_{1},n_{2},n_{3})\in T_{2}$, then $\operatorname{uc}(S,\mathscr{O}_{S}(1))=1$.
-The construction of Ulrich rank-two vector bundles to prove Theorem 1.2 uses the Cayley–Bacharach–Serre correspondence method (cf.
-Theorem 5.1).
-Additionally, we compute the Ulrich complexity in some low-degree cases, completing the picture for even bidouble planes (cf.
-Proposition 4.6).
-In [4, Theorem 1.1], Basu and Pal establish the existence of rank two Ulrich bundles on certain surfaces of general type.
-However, our results lie outside the scope of their framework, as they assume that the surfaces are polarized by their canonical bundle.
-This is not our situation, since we investigate Ulrich bundles with respect to a non–very ample divisor.
-Moreover, [4, Theorem 1.1] requires that the canonical class generate the Picard group.
 The paper is organized as follows.
@@
 On the positive side, we also construct explicit Ulrich line bundles in the cases $(n_{1},n_{2},n_{3})=(0,2,2)$ and $(0,2,4)$.
-Finally, in Section 5, we prove Theorem 1.2, showing that the Ulrich complexity of an even bidouble plane with sufficiently large degree is equal to $2$.
 Remark 1.3.
@@
     $h^{0}(S,\mathscr{E}(-H))=h^{0}(S,\mathscr{E}^{\vee}(2H+K_{S}))=0$ and Equalities (1) and (2) hold.
-Definition 2.4.
-Let $X$ be a smooth projective variety of dimension $d$, and let $H$ be an ample and globally generated divisor on $X$.
-A rank $2$ vector bundle $\mathscr{E}$ on $X$ is called special Ulrich if it is Ulrich with respect to $H$ and $c_{1}(\mathscr{E})=(d+1)H+K_{X}$.
-Corollary 2.5 ([12, Proposition 5.1.3]).
-Let $(S,H)$ be a smooth projective polarized surface and let $\mathscr{E}$ be a rank $2$ vector bundle on $S$. Then $\mathscr{E}$ is a special Ulrich bundle if and only if it satisfies
-  1. (1)
-    $c_{1}(\mathscr{E})\equiv 3H+K_{S},$
-  2. (2)
-    $c_{2}(\mathscr{E})=\frac{1}{2}(5H^{2}+3H\cdot K_{S})+2\chi(\mathscr{O}_{S}),$
-  3. (3)
-    $h^{0}(S,\mathscr{E})>0,$ and
-  4. (4)
-    $h^{0}(S,\mathscr{E}(-H))=0.$
-A vector bundle is called initialized with respect to $H$ if it satisfies the conditions (3) and (4).
-A fruitful approach for constructing rank 2 vector bundles on surfaces with prescribed Chern classes is provided by Cayley–Bacharach–type results.
-For this reason, we state the following theorem, which will play a central role in the proof of Theorem 1.2.
-Theorem 2.6 ([26, Theorem 7]).
-Let $X$ be a smooth, projective variety of dimension at least $2$, and let $Z\subset X$ be a subscheme of pure codimension $2$. The following statements are equivalent:
-  1. (1)
-    $Z$ is the zero subscheme of a section of a rank $2$ vector bundle $\mathscr{E}$,
-  2. (2)
-    there exist three hypersurfaces $E_{1},\,E_{2}$ and $E_{3}$ such that $E_{1}$ and $E_{2}$ have no common components, $Z=E_{1}\cap E_{2}\setminus(E_{1}\cap E_{2}\cap E_{3})$, and $E_{1}\cap E_{2}\cap E_{3}$ is a pure codimension $2$ and Cohen–Maculay.
-If (1) or (2) holds, then $\mathscr{E}$ fits into a short exact sequence
-$0\longrightarrow\mathscr{O}_{X}\longrightarrow\mathscr{E}\longrightarrow\mathscr{I}_{Z}(E_{1}+E_{2}-E_{3})\longrightarrow 0,$ (3)
-which implies $c_{1}(\mathscr{E})=E_{1}+E_{2}-E_{3}$ and $c_{2}({\mathscr{E}})=\#Z$.
-Remark 2.7.
-The sequence (3) can be found at the end of the proof of [26, Theorem 7].
 2.2. Bidouble covers of surfaces
@@
 This observation suggests that the existence of Ulrich line bundles on $S$ in these cases is unlikely, as illustrated by Proposition 4.4.
-5. rank two Ulrich bundles on general even bidouble planes
-In this section, we prove Theorem 1.2.
-We show that, in contrast to the situation for Ulrich line bundles, special Ulrich bundles exist for almost every even generic bidouble cover $S$ of the projective plane.
-Recall that $H$ denotes the polarization $\pi^{*}\mathscr{O}_{{\mathbb{P}}^{2}}(1)$.
-By Corollary 2.5, in order to construct a special Ulrich bundle with respect to $H$, it suffices to produce a rank-two vector bundle $\mathscr{E}$ satisfying
-$c_{1}(\mathscr{E})\equiv 3H+K_{S},\qquad c_{2}(\mathscr{E})=\frac{1}{2}(5H^{2}+3H\cdot K_{S})+2\chi(\mathscr{O}_{S}),$
-together with the cohomological conditions
-$h^{0}(S,\mathscr{E})>0,\qquad h^{0}(S,\mathscr{E}(-H))=0.$
-The key idea is to apply Theorem 2.6 to construct a vector bundle of rank two satisfying these requirements. More precisely, we choose hypersurfaces $E_{1}$, $E_{2}$, $E_{3}$ such that
-$3H+K_{S}\equiv E_{1}+E_{2}-E_{3},\qquad\text{and}\qquad\#Z=\frac{1}{2}(5H^{2}+3H\cdot K_{S})+2\chi(\mathscr{O}_{S}),$
-where $Z=(E_{1}\cap E_{2})\setminus(E_{1}\cap E_{2}\cap E_{3})$.
-According to Theorem 2.6, these data produce a rank-two vector bundle $\mathscr{E}$ with the desired Chern classes.
-The remainder of the proof consists in using the short exact sequence (3) to verify the required cohomology vanishings.
-Theorem 5.1 (Ulrich bundles of rank two).
-Let $\pi\colon S\to\mathbb{P}^{2}$ be an even smooth bidouble plane branched along smooth curves $D_{1},D_{2},D_{3}$ of degrees $n_{1},n_{2},n_{3}$, respectively.
-If no permutation of $(n_{1},n_{2},n_{3})$ is equal to $(0,2,2)$, then $S$ carries a special Ulrich bundle $\mathscr{E}$ with respect to $H$.
-Proof.
-Let $_{i}=n_{i}/2, m=m$$m_{1}+m_{2}+m_{3}$ and $M=m^{2}+m_{1}^{2}+m_{2}^{2}+m_{3}^{2}$, and notice that $M$ is even.
-By Corollary 2.5, a rank-two vector bundle $\mathscr{E}$ is a special Ulrich bundle with respect to $H$ if and only if it satisfies
-$c_{1}(\mathscr{E})\equiv mH,\qquad c_{2}(\mathscr{E})=M,\qquad h^{0}(S,\mathscr{E})>0,\qquad\text{and}\qquad h^{0}(S,\mathscr{E}(-H))=0.$
-Now, we define the hypersurfaces $E_{1}$, $E_{2}$, and $E_{3}$ according to the reduction of $M$ modulo $4$.
-If $M\equiv 0\pmod{4}$, set $E_{1}\colon=\pi^{*}(\ell)$, $E_{2}\colon=\pi^{*}(C)$, and $E_{3}\colon=\pi^{*}(C^{\prime})$, where $\ell$ is a general line, and $C$ and $C^{\prime}$ are smooth plane curves in general position of degrees $M/4$, and M/4+1mM/4+1-m, respectively.
-On the other hand, if $M\equiv 2\pmod{4}$, set $E_{2}\colon=\pi^{*}(C)$, and $E_{3}\colon=\pi^{*}(C^{\prime})$, where $C$ and $C^{\prime}$ are smooth plane curves in general position of degrees $(M+2)/4$, and (M+2)/4+1m(M+2)/4+1-m, respectively, and $E_{1}\colon=\pi^{*}(\ell)$, where $\ell$ is a line that intersects exactly one point of $C$ and the $D_{i}$’s.
-In any case, since $m\geq 3$, $\deg(C)\geq m$, in particular, the degree of $C^{\prime}$ is positive. Due to the generality, we have that $E_{1},\,E_{2}$, and $E_{3}$ have no common point.
-Let $Z\subset S$ be the zero-dimensional subscheme
-$Z\coloneqq E_{1}\cap E_{2}\;\setminus\;(E_{1}\cap E_{2}\cap E_{3}).$
-Since the triple intersection is empty and because of the choice of $E_{i}$’s, if $M\equiv 0\pmod{4}$, we have $Z=4\left(\frac{M}{4}\right)=M$.
-In the case $M\equiv 2\pmod{4}$, there is a single point in $\ell\cap C$ which if ramified via $\pi$, then $Z=4\left(\frac{M+2}{4}-1\right)+2=M$.
-By construction, the data $(E_{1},E_{2},E_{3},Z)$ satisfy Theorem 2.6 (2).
-Hence, there exists a rank two vector bundle $\mathscr{E}$ fitting into
-$0\longrightarrow\mathscr{O}_{S}\longrightarrow\mathscr{E}\longrightarrow\mathscr{I}_{Z}(E_{1}+E_{2}-E_{3})\longrightarrow 0.$ (13)
-Since $E_{1}+E_{2}-E_{3}\equiv mH$, we obtain $c_{1}(\mathscr{E})\equiv mH$, and $c_{2}(\mathscr{E})=\#Z=M$.
-Now, we claim that $h^{0}(S,\mathscr{E}(-H))=0$. Twisting (13) by $\mathscr{O}_{S}(-H)$ gives
-$0\longrightarrow\mathscr{O}_{S}(-H)\longrightarrow\mathscr{E}(-H)\longrightarrow\mathscr{I}_{Z}((m-1)H)\longrightarrow 0.$ (14)
-Since $h^{0}(S,{\mathscr{O}}_{S}(-H))=0$, it is enough to prove that $h^{0}(S,\mathscr{I}_{Z}((m-1)H))=0$.
-For the sake of contradiction, assume that there exists a nonzero section of $\mathscr{I}_{Z}((m-1)H)$, and let $D\in|(m-1)H|$ be the corresponding effective divisor.
-Then $Z\subset D$.
-Intersecting with $E_{1}\in|H|$, we obtain $D\cdot E_{1}=4(m-1)$.
-On the other hand, since $Z\subset E_{1}$ and every point of $Z$ contributes at least one to the intersection multiplicity, we have $D\cdot E_{1}\;\geq\;\#Z=M$.
-Because $m_{1}^{2}+m_{2}^{2}+m_{3}^{2}\geq\frac{1}{3}(m_{1}+m_{2}+m_{3})^{2}=\frac{1}{3}m^{2}$, we obtain $M\geq 4m^{2}/3$.
-For $m\geq 3$, this gives $M>4(m-1)=D\cdot E_{1}$, a contradiction.
-Therefore $h^{0}\bigl(S,\mathscr{I}_{Z}((m-1)H)\bigr)=0$, and hence $h^{0}(S,\mathscr{E}(-H))=0$.
-Finally, the inclusion $\mathscr{O}_{S}\hookrightarrow\mathscr{E}$ induces a nonzero section $s\in H^{0}(S,\mathscr{E})$. Combining all the above, the vector bundle $\mathscr{E}$ satisfies
-$c_{1}(\mathscr{E})\equiv mH,\qquad c_{2}(\mathscr{E})=M,\qquad h^{0}(S,\mathscr{E})>0,\qquad\text{and}\qquad h^{0}(S,\mathscr{E}(-H))=0,$
-as desired. ∎
-Proof of Theorem 1.2.
-Let us recall the definition of $T_{1}$ and $T_{2}$:
-$T_{1}=\{(0,4,2n)\mid n\geq 2\}\;\cup\;\{(2,2,2n)\mid n\geq 1\},\qquad\text{and}\qquad T_{2}=\{(0,2,2),\,(0,2,4)\}.$
-To begin with, assume that $S$ is an odd bidouble cover. Theorem 4.5 implies that $S$ does not admit an Ulrich line bundle with respect to $H$. Consequently, $\operatorname{uc}(S,H)>1$.
-Now, assume that $S$ is an even bidouble cover.
-First, Suppose that $(n_{1},n_{2},n_{3})\notin T_{1}\cup T_{2}$.
-Theorem 4.5 implies that $S$ does not admit an Ulrich line bundle with respect to $H$.
-On the other hand, Theorem 5.1 implies the existence of an Ulrich rank two vector bundle.
-Consequently, $\operatorname{uc}(S,H)=2$.
-Now, if $(n_{1},n_{2},n_{3})\in T_{1}$ Theorem 5.1 implies that $\operatorname{uc}(S,H)\leq 2$.
-Finally, if $(n_{1},n_{2},n_{3})\in T_{2}$ Proposition 4.6 implies that $S$ admits an Ulrich line bundle with respect to $H$, which yields $\operatorname{uc}(S,H)=1$.
-∎